# Optimal High-Probability Regret for Online Convex Optimization with Two-Point Bandit Feedback

### 26 Mar 2026 v1

@@ Abstract
We consider the problem of Online Convex Optimization (OCO) with two-point bandit feedback in an adversarial environment. In this setting, a player attempts to minimize a sequence of adversarially generated convex loss functions, while only observing the value of each function at two points. While it is well-known that two-point feedback allows for gradient estimation, achieving tight high-probability regret bounds for strongly convex functions still remained open as highlighted by Agarwal et al. (2010). The primary challenge lies in the heavy-tailed nature of bandit gradient estimators, which makes standard concentration analysis difficult. In this paper, we resolve this open challenge by providing the first high-probability regret bound of $O(d(\log T+\log(1/\delta))/\mu)$ for $\mu$-strongly convex losses. Our result is minimax optimal with respect to both the time horizon $T$ and the dimension $d$.
@@ 1 Introduction
To bridge the gap between bandit and full-information feedback where the gradient information of the loss function $\ell_{t}(\cdot)$ can be obtained, Agarwal et al. (2010) introduced the multi-point bandit setting. In their seminal work, they demonstrated that evaluating the loss function at two randomized points allows the player to construct gradient estimators that mimic the behavior of true gradients. However, as explicitly discussed in their paper, achieving tight high-probability regret bounds for strongly convex loss functions under two-point feedback remained an unresolved challenge. Specifically, while expectation bounds were established, the heavy-tailed nature of bandit estimators against an adversary has historically made high-confidence guarantees elusive.
@@
This regret is minimax optimal both in $T$ (Abernethy et al., 2009) and dimension $d$ (Duchi et al., 2015).
@@
Furthermore, our result significantly improves the dependency on the dimension $d$. We reduce the regret from the $O(d^{2})$ factor found in prior work (Agarwal et al., 2010). This improvement is achieved by extending the sophisticated geometric and probabilistic techniques introduced by Shamir (2017), which were used to prove a regret holding in expectation, to the high-probability regime. By generalizing these techniques to accommodate high-confidence concentration, we match the information-theoretic lower bounds for this online convex optimization setting.
@@ 2.1 Notation
We first define our notation will appear in various parts of the analysis. Let $\|\cdot\|$ denote the Euclidean norm in $\mathbb{R}^{d}$, and let $\mathcal{B}=\{\bm{x}\in\mathbb{R}^{d}:\|\bm{x}\|\leq 1\}$ be the unit ball centered at the origin. We assume that the convex set $\mathcal{K}$ is compact and has a nonempty interior. If not, we can always map $\mathcal{K}$ to a lower dimensional space. More specifically, we assume that there exist $D>0$ such that
$\mathcal{K}\subseteq D\mathcal{B}.$ (1)
@@ 2.2 Preliminaries
Lemma 1 (Nesterov and Spokoiny (2017))
@@
Lemma 2 (Lemma 10 of Shamir (2017))
@@
Lemma 3 (Lemma 2 of Agarwal et al. (2010))
@@ 3 Main Results
Theorem 4
Suppose the loss function $\ell_{t}(\cdot)$ is $\mu$-strongly convex and $G$-Lipschitz continuous. The constraint set $\mathcal{K}$ is convex and satisfies Eq. (1). Let the sequence $\{\bm{x}_{t}\}$ be generated by Algorithm 1. Given $0<\delta<1$, by setting the step size $\eta_{t}=\frac{2}{\mu t}$, $\alpha=\frac{dG\log\frac{1}{\delta}}{T\mu}$ and $\xi=(d+1)\log\frac{1}{\delta}/T$, then with a probability at least $1-\delta$, it holds that for any fixed $\bm{x}\in\mathcal{K}$
@@
$\leq$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{(32c_{0}+3)dG^{2}\log\frac{1}{\delta}}{\mu}+\frac{7(d+1)GD\log\frac{1}{\delta}}{3},$
@@
Above regret matches the one of Agarwal et al. (2010) which holds in expectation except that our regret bound has a linear dependence on the dimension while Agarwal et al. (2010) quadratically depends on the dimension. Thus, our result resolves the open problem left by Agarwal et al. (2010) whether one can achieve a $O(\log T)$ regret with high-probability for strongly convex loss functions under two-point feedback. Furthermore, the regret in Eq. (6) is the first one which achieves the minimax optimal both in $T$ (Abernethy et al., 2009) and dimension $d$ (Duchi et al., 2015).
@@ 3.1 Detailed Proofs
Our proof will revisit the proof of online convex optimization which achieves the $O(\log T)$ regret for the strongly convex function (Hazan et al., 2007). This is the key step to achieve our two-point feed-back algorithm with $O(\log T)$ regret holding in a high probability.
Lemma 5
@@
Proof First, by the property of projection and the convexity of $\mathcal{K}$, we have
$\|\bm{x}_{t+1}-\bm{x}\|^{2}=$ $\left\|\Pi_{(1-\xi)\mathcal{K}}(\bm{x}_{t}-\eta_{t}\bm{g}_{t})\right\|^{2}\leq\left\|\bm{x}_{t}-\eta_{t}\bm{g}_{t}\right\|^{2}$
@@
Then, by setting step size $\eta_{t}=\frac{1}{\mu t}$ and upper bounding $\mathbb{E}[\|\bm{g}_{t}\|^{2}]\leq O(d^{2})$, then following the proof of Hazan et al. (2007), we can obtain the regret $O(d^{2}\log T)$ in expectation achieved in Agarwal et al. (2010).
To achieve a high-probability probability, one has to bound the extra term $\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$ while it will disappear if we only need a result holding in expectation. However, it will incur an additional term of $O(\sqrt{T})$ by directly using the Hoffding-Azuma’s inequality (Agarwal et al., 2010).
Next, the following lemma gives a novel upper bound of $\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$ holding with a high probability by using the Freedman’s inequality. This is a key step to overcome the additional term of $O(\sqrt{T})$ caused by directly using the Hoffding-Azuma’s inequality.
Lemma 6
Suppose the loss function $\ell_{t}(\bm{x})$ is $G$-Lipschitz continuous and the convex set $\mathcal{K}$ satisfies Eq. (1). Given $0<\delta<1$, then with a probability at least $1-\delta$, it holds that
$\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle\leq\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu},$ (9)
where $c_{0}$ is an absolute constant.
Proof Let us denote
$Z_{t}:=\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$
By Lemma 2, we can obtain that
$\mathbb{E}[Z_{t}\mid\bm{x}_{t}]=0.$
Furthermore,
$|Z_{t}|\leq$ $\|\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t}\|\cdot\|\bm{x}_{t}-\bm{x}\|\leq(\|\nabla\hat{\ell}_{t}(\bm{x}_{t})\|+\|\bm{g}_{t}\|)\cdot\|\bm{x}_{t}-\bm{x}\|$
$\leq$ $(G+dG)\cdot 2D=2(d+1)GD,$
where the last inequality is because the loss function $\ell_{t}(\bm{x})$ is $G$-Lipschitz continuous, Lemma 1, Eq. (1), and
@@
Next, we have
$\mathbb{E}[Z_{t}^{2}\mid\bm{x}_{t}]=$ $\mathbb{E}[\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}]$
$=$ $\mathbb{E}\left[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}\right]+\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle^{2}$
$-2\mathbb{E}\left[\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle\cdot\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle\mid\bm{x}_{t}\right]$
$=$ $\mathbb{E}\left[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}\right]-\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\rangle^{2}$
$\leq$ $\mathbb{E}\left[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\mid\bm{x}_{t}\right]\leq\mathbb{E}[\|\bm{g}_{t}\|^{2}]\cdot\|\bm{x}_{t}-\bm{x}\|^{2}$
$\leq$ $c_{0}dG^{2}\|\bm{x}_{t}-\bm{x}\|^{2},$
where the third equality and the last inequality are both because of Lemma 2.
Thus, we have
$\sum_{t=1}^{T}\mathbb{E}[Z_{t}^{2}\mid\bm{x}_{t}]\leq c_{0}dG^{2}\cdot\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}.$
Then, by Lemma 11 (Freedman’s inequality) with $R=2(d+1)GD$ and $\sigma^{2}=c_{0}dG^{2}\cdot\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}$, we can obtain that with a probability at least $1-\delta$, it holds that
$\sum_{t=1}^{T}Z_{t}\leq$ $\frac{2R\log\frac{1}{\delta}}{3}+\sqrt{2\sigma^{2}\log\frac{1}{\delta}}\leq\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\sqrt{2c_{0}dG^{2}\log\frac{1}{\delta}\cdot\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}}$
$\leq$ $\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$
@@
With a properly chosen step size $\eta_{t}=2/(\mu t)$ instead of $1/(\mu t)$ chosen in Agarwal et al. (2010) and Hazan et al. (2007), we can obtain the following regret where no extra $\sqrt{T}$ term shows.
Lemma 7
@@
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$ (10)
Proof We have
@@
$+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}-\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}$
@@
$+\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu},$
@@
Combining above results, we can obtain the final result.
Finally, we will show $\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}$ is upper bounded by $O(d\log T)$. This step is the key step that guarantees Algorithm 1 can achieve a regret linear dependence on the dimension. This is achieved by extending the technique introduced by Shamir (2017). The technique in Shamir (2017) only proves the regret holding in expectation. In this paper, we extend it to achieve the regret holding with a high probability.
Lemma 8
@@
Proof We have
@@
$\leq$ $\frac{d^{2}}{\alpha^{2}\mu}\cdot K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right)$
$=$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right),$
where the last inequality is because of Lemma 19.
Lemma 9
@@
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$ (12)
Proof Combining Eq. (10) and Eq. (11), we can obtain that
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq$ $\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}$
$\leq$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{4(d+1)GD\log\frac{1}{\delta}}{3}+\frac{32c_{0}dG^{2}\log\frac{1}{\delta}}{\mu}.$
@@
Proof [Proof of Theorem 4] Combining Lemma 3 and Lemma 9, we can obtain that
@@
$=$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{(32c_{0}+3)dG^{2}\log\frac{1}{\delta}}{\mu}+\frac{7(d+1)GD\log\frac{1}{\delta}}{3},$
where the second inequality is also because $(1-\xi)\bm{x}\in(1-\xi)\mathcal{K}$ if $\bm{x}\in\mathcal{K}$ and the last inequality is because of $\alpha=\frac{dG\log\frac{1}{\delta}}{T\mu}$ and $\xi=(d+1)\log\frac{1}{\delta}/T$.

### 05 Apr 2026 v3

@@ Abstract
We consider the problem of Online Convex Optimization (OCO) with two-point bandit feedback. In this setting, a player attempts to minimize a sequence of adversarially generated convex loss functions, while only observing the value of each function at two points. While it is well-known that two-point feedback allows for gradient estimation, achieving tight high-probability regret bounds for strongly convex functions still remained open as highlighted by Agarwal et al., (2010). The primary challenge lies in the heavy-tailed nature of bandit gradient estimators, which makes standard concentration analysis difficult. In this paper, we resolve this open challenge and provide the first high-probability regret bound of $O(d(\log T+\log(1/\delta))/\mu)$ for $\mu$-strongly convex losses. Our result is minimax optimal with respect to both the time horizon $T$ and the dimension $d$.
@@ 1 Introduction
To bridge the gap between bandit and full-information feedback where the gradient information of the loss function $\ell_{t}(\cdot)$ can be obtained, Agarwal et al., (2010) introduced the multi-point bandit setting. In their seminal work, they demonstrated that evaluating the loss function at two randomized points allows the player to construct gradient estimators that mimic the behavior of true gradients. However, as explicitly discussed in their paper, achieving tight high-probability regret bounds for strongly convex loss functions under two-point feedback remained an unresolved challenge. Specifically, while expectation bounds were established, the heavy-tailed nature of bandit estimators against an adversary has historically made high-confidence guarantees elusive.
@@
This regret is minimax optimal in the time horizon $T$ (Abernethy et al.,, 2009) and the dimension $d$ of the problem (Duchi et al.,, 2015).
@@
Furthermore, our result significantly improves the horizon-dependent term. We reduce the $O(d^{2})$ factor in the horizon-dependent term found in prior work (Agarwal et al.,, 2010) to the linear dependence. This improvement is achieved by extending the sophisticated geometric and probabilistic techniques introduced by Shamir, (2017), which were used to prove a regret holding in expectation, to the high-probability regime. By generalizing these techniques to accommodate high-confidence concentration, we match the information-theoretic lower bounds for this online convex optimization setting.
@@ 2.1 Notation
We first define our notation will appear in various parts of the analysis. Let $\|\cdot\|$ denote the Euclidean norm in $\mathbb{R}^{d}$, and let $\mathcal{B}=\{\bm{x}\in\mathbb{R}^{d}:\|\bm{x}\|\leq 1\}$ be the unit ball centered at the origin. We assume that the convex set $\mathcal{K}$ is compact and has a nonempty interior. If not, we can always map $\mathcal{K}$ to a lower dimensional space. More specifically, we assume that there exist $r,D>0$ such that
$r\mathcal{B}\subseteq\mathcal{K}\subseteq D\mathcal{B}.$ (1)
@@ 2.2 Preliminaries
Lemma 1 (Nesterov & Spokoiny, (2017)).
@@
Lemma 2 (Lemma 10 of Shamir, (2017)).
@@
Lemma 3 (Lemma 2 of Agarwal et al., (2010)).
@@ 3 Main Results
Theorem 1.
Suppose the loss function $\ell_{t}(\cdot)$ is $\mu$-strongly convex and $G$-Lipschitz continuous. The constraint set $\mathcal{K}$ is convex and satisfies Eq. (1). Let the sequence $\{\bm{x}_{t}\}$ be generated by Algorithm 1. Given $0<\delta<1$, by setting the step size $\eta_{t}=\frac{2}{\mu t}$, $\alpha=\frac{\log T}{T}$ and $\xi=\alpha/r$, then with a probability at least $1-\delta$, it holds that for any fixed $\bm{x}\in\mathcal{K}$
@@
$\leq$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+2(d+1)GD\log\frac{1}{\delta}+G\log T\left(3+\frac{D}{r}\right),$
@@
Above regret matches the one of Agarwal et al., (2010) which holds in expectation except that our regret bound has a linear dimension dependence in the horizon-dependent term while Agarwal et al., (2010) quadratically depends on the dimension. Thus, our result resolves the open problem left by Agarwal et al., (2010) whether one can achieve a $O(\log T)$ regret with high-probability for strongly convex loss functions under two-point feedback. Furthermore, the regret in Eq. (6) is the first one which achieves the minimax optimal both in $T$ (Abernethy et al.,, 2009) and dimension $d$ (Duchi et al.,, 2015).
@@ 3.1 Detailed Proofs
Our proof will revisit the proof of online convex optimization which achieves the $O(\log T)$ regret for the strongly convex function (Hazan et al.,, 2007). This is the key step to achieve our two-point feed-back algorithm with $O(\log T)$ regret holding in a high probability.
Lemma 4.
@@
Proof.
First, by the property of projection and the convexity of $\mathcal{K}$, we have
$\|\bm{x}_{t+1}-\bm{x}\|^{2}=$ $\left\|\Pi_{(1-\xi)\mathcal{K}}(\bm{x}_{t}-\eta_{t}\bm{g}_{t})-\bm{x}\right\|^{2}\leq\left\|\bm{x}_{t}-\eta_{t}\bm{g}_{t}-\bm{x}\right\|^{2}$
@@
Then, by setting step size $\eta_{t}=\frac{1}{\mu t}$ and upper bounding $\mathbb{E}[\|\bm{g}_{t}\|^{2}]\leq O(d^{2})$, then following the proof of Hazan et al., (2007), we can obtain the regret $O(d^{2}\log T)$ in expectation achieved in Agarwal et al., (2010).
To achieve a high-probability probability, one has to bound the extra term $\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$ while it will disappear if we only need a result holding in expectation. However, it will incur an additional term of $O(\sqrt{T})$ by directly using the Hoffding-Azuma’s inequality (Agarwal et al.,, 2010).
Next, the following lemma gives a high-probability upper bound on $\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle$. The proof uses a Freedman-style exponential supermartingale, rather than Freedman’s tail inequality with a random variance threshold.
Lemma 5.
Suppose the loss function $\ell_{t}(\bm{x})$ is $G$-Lipschitz continuous and the convex set $\mathcal{K}$ satisfies Eq. (1). Given $0<\delta<1$, then with probability at least $1-\delta$, it holds that
$\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle\leq\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{4dG^{2}}{\mu}\log\frac{1}{\delta}+2(d+1)GD\log\frac{1}{\delta}.$ (9)
Proof.
Let
$Z_{t}:=\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle,\qquad\mathcal{F}_{t}:=\sigma(\bm{u}_{1},\ldots,\bm{u}_{t}).$
Since $\bm{x}_{t}$ is measurable with respect to $\mathcal{F}_{t-1}$ and Lemma 2 gives
$\mathbb{E}[\bm{g}_{t}\mid\bm{x}_{t}]=\nabla\hat{\ell}_{t}(\bm{x}_{t}),$
we have
$\mathbb{E}[Z_{t}\mid\mathcal{F}_{t-1}]=0.$
Hence $(Z_{t})_{t=1}^{T}$ is a martingale difference sequence.
Next, we bound $|Z_{t}|$. By Cauchy–Schwarz,
$|Z_{t}|$ $\leq\|\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t}\|\cdot\|\bm{x}_{t}-\bm{x}\|$
$\leq\bigl(\|\nabla\hat{\ell}_{t}(\bm{x}_{t})\|+\|\bm{g}_{t}\|\bigr)\|\bm{x}_{t}-\bm{x}\|.$
Because $\ell_{t}$ is $G$-Lipschitz continuous, Lemma 1 implies $\|\nabla\hat{\ell}_{t}(\bm{x}_{t})\|\leq G$. Moreover,
@@
Using Eq. (1), we have $\|\bm{x}_{t}-\bm{x}\|\leq 2D$, and therefore
$|Z_{t}|\leq 2(d+1)GD.$
Set
$b:=2(d+1)GD.$
Now we bound the conditional second moment. Since $\mathbb{E}[\bm{g}_{t}\mid\mathcal{F}_{t-1}]=\nabla\hat{\ell}_{t}(\bm{x}_{t})$, we have
$\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}]$ $=\mathbb{E}\Big[\big\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\big\rangle^{2}\,\Big|\,\mathcal{F}_{t-1}\Big]$
$=\mathbb{E}\Big[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\,\Big|\,\mathcal{F}_{t-1}\Big]-\big\langle\nabla\hat{\ell}_{t}(\bm{x}_{t}),\bm{x}_{t}-\bm{x}\big\rangle^{2}$
$\leq\mathbb{E}\Big[\langle\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle^{2}\,\Big|\,\mathcal{F}_{t-1}\Big].$
Let $\bm{y}_{t}:=\bm{x}_{t}-\bm{x}$, which is $\mathcal{F}_{t-1}$-measurable. Then
$\langle\bm{g}_{t},\bm{y}_{t}\rangle^{2}$ $=\left(\frac{d}{2\alpha}\big(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})\big)\langle\bm{u}_{t},\bm{y}_{t}\rangle\right)^{2}$
$=\frac{d^{2}}{4\alpha^{2}}\big(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\ell_{t}(\bm{x}_{t}-\alpha\bm{u}_{t})\big)^{2}\langle\bm{u}_{t},\bm{y}_{t}\rangle^{2}$
$\leq\frac{d^{2}}{4\alpha^{2}}\cdot(2\alpha G)^{2}\langle\bm{u}_{t},\bm{y}_{t}\rangle^{2}$
$=d^{2}G^{2}\langle\bm{u}_{t},\bm{y}_{t}\rangle^{2},$
where we used the $G$-Lipschitz continuity of $\ell_{t}$. Taking conditional expectation and using the standard identity $\mathbb{E}[\langle\bm{u}_{t},\bm{y}_{t}\rangle^{2}\mid\mathcal{F}_{t-1}]=\|\bm{y}_{t}\|^{2}/d$, we get
$\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}]\leq dG^{2}\|\bm{x}_{t}-\bm{x}\|^{2}.$
Define the predictable quadratic variation
$V_{T}:=\sum_{t=1}^{T}\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}].$
Then
$V_{T}\leq dG^{2}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}.$
We now apply the Freedman-style exponential supermartingale lemma (Lemma 10) to $(Z_{t})_{t=1}^{T}$: for every $\lambda\in[0,3/b)$, with probability at least $1-\delta$,
$\sum_{t=1}^{T}Z_{t}\leq\frac{\lambda}{2(1-\lambda b/3)}V_{T}+\frac{1}{\lambda}\log\frac{1}{\delta}.$
Substituting the bound on $V_{T}$ yields
$\sum_{t=1}^{T}Z_{t}\leq\frac{\lambda dG^{2}}{2(1-\lambda b/3)}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{1}{\lambda}\log\frac{1}{\delta}.$
Choose
$\lambda:=\frac{1}{b+4dG^{2}/\mu}.$
Then $\lambda<1/b<3/b$, so the lemma applies, and
$1-\lambda b/3\geq 1-\lambda b\geq 0.$
More usefully, since $\lambda b\leq 1$, we have
$1-\lambda b/3\geq\frac{2}{3},$
and therefore
$\frac{\lambda dG^{2}}{2(1-\lambda b/3)}\leq\frac{3}{4}\lambda dG^{2}\leq\frac{\mu}{4},$
because $\lambda\leq\mu/(3dG^{2})$ under the chosen value. Hence
$\sum_{t=1}^{T}Z_{t}\leq\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\left(b+\frac{4dG^{2}}{\mu}\right)\log\frac{1}{\delta}.$
Recalling $b=2(d+1)GD$, we conclude that with probability at least $1-\delta$,
$\sum_{t=1}^{T}\langle\nabla\hat{\ell}_{t}(\bm{x}_{t})-\bm{g}_{t},\bm{x}_{t}-\bm{x}\rangle\leq\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{4dG^{2}}{\mu}\log\frac{1}{\delta}+2(d+1)GD\log\frac{1}{\delta}.$
This completes the proof. ∎
@@
With a properly chosen step size $\eta_{t}=2/(\mu t)$ instead of $1/(\mu t)$ chosen in Agarwal et al., (2010) and Hazan et al., (2007), we can obtain the following regret where no extra $\sqrt{T}$ term shows.
Lemma 6.
@@
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4dG^{2}}{\mu}\log\frac{1}{\delta}+2(d+1)GD\log\frac{1}{\delta}.$ (10)
Proof.
We have
@@
$-\frac{\mu}{4}\sum_{t=1}^{T}\|\bm{x}_{t}-\bm{x}\|^{2}+\frac{4dG^{2}}{\mu}\log\frac{1}{\delta}+2(d+1)GD\log\frac{1}{\delta}$
@@
$+\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4dG^{2}}{\mu}\log\frac{1}{\delta}+2(d+1)GD\log\frac{1}{\delta},$
@@
Combining above results, we can obtain the final result. ∎
Finally, we will show $\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}$ is upper bounded by $O(d\log T)$. This step is the key step that guarantees Algorithm 1 can achieve a regret linear dependence on the dimension. This is achieved by extending the technique introduced by Shamir, (2017). The technique in Shamir, (2017) only proves the regret holding in expectation. In this paper, we extend it to achieve the regret holding with a high probability.
Lemma 7.
@@
Proof.
First, we have
$\sum_{t=1}^{T}\frac{\left(\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]\right)^{2}}{t}\leq K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right),$
where the last inequality is because of Lemma 17.
Furthermore, by the symmetry of spherical distribution, the random variables $\mathbb{E}_{\bm{u}}\!\left[\ell_{t}(\bm{x}_{t}-\alpha\bm{u})\right]-\ell_{t}(\bm{x}_{t}-\alpha\bm{u})$ and $\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})-\mathbb{E}_{\bm{u}_{t}}[\ell_{t}(\bm{x}_{t}+\alpha\bm{u}_{t})]$ share the same distribution. Thus, it also holds that
$\sum_{t=1}^{T}\frac{\left(\mathbb{E}_{\bm{u}}\!\left[\ell_{t}(\bm{x}_{t}-\alpha\bm{u})\right]-\ell_{t}(\bm{x}_{t}-\alpha\bm{u})\right)^{2}}{t}\leq K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right).$
Combining with Eq. (18), we can obtain that We have
@@
$\leq$ $\frac{3d^{2}}{4\alpha^{2}\mu}\cdot\left(K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right)+K\cdot\frac{(G\alpha)^{2}}{d}\cdot\left(\log T+\log\frac{1}{\delta}\right)\right)$
$=$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right).$
Lemma 8.
@@
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+2(d+1)GD\log\frac{1}{\delta}.$ (12)
Proof.
Combining Eq. (10) and Eq. (11), we can obtain that
$\sum_{t=1}^{T}\hat{\ell}_{t}(\bm{x}_{t})-\hat{\ell}_{t}(\bm{x})\leq$ $\sum_{t=1}^{T}\frac{\left\|\bm{g}_{t}\right\|^{2}}{\mu t}+\frac{4dG^{2}}{\mu}\log\frac{1}{\delta}+2(d+1)GD\log\frac{1}{\delta}$
$\leq$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+\frac{4dG^{2}}{\mu}\log\frac{1}{\delta}+2(d+1)GD\log\frac{1}{\delta}$
$=$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+2(d+1)GD\log\frac{1}{\delta}.$
@@
Proof of Theorem 1.
Combining Lemma 3 and Lemma 8, we can obtain that
@@
$=$ $K\cdot\frac{dG^{2}}{\mu}\cdot\left(\log T+\log\frac{1}{\delta}\right)+2(d+1)GD\log\frac{1}{\delta}+G\log T\left(3+\frac{D}{r}\right),$
where the second inequality is also because $(1-\xi)\bm{x}\in(1-\xi)\mathcal{K}$ if $\bm{x}\in\mathcal{K}$ and the last inequality is because of $\alpha=\frac{\log T}{T}$ and $\xi=\alpha/r$. ∎